Home Clean room MedAustron HEPHY testbeams old
testbeam-DESY-Oct2017 testbeam-DESY-April2017 testbeam-SPS2015 testbeam-SPS2014 testbeam-DESY14 testbeam-SPS12 testbeam-SPS11 testbeam-SPS10 testbeam-SPS09 testbeam-SPS08 LP-TPC
  HEPHY logbook for testbeam at SPS September/October 2010  Not logged in ELOG logo
Message ID: 102     Entry time: Fri Oct 8 23:43:07 2010
Author: Christian Irmler 
Project: BELLE 
Subject: run 
Run Number: run055 
Events: 12316 
StartTime: 08.10.2010 23:42:51 
EndTime: 09.10.2010 00:14:56  

Belle stack w/o Origami2 after irradiation
HV=80V / 255 µA
cooling: Origami only, chiller set to -25°C,
2 suspicious APVs on n side of Micron due to burnt strip

same position as in previous run

telescope run 20197  - wrong config file @ telescope -> new run with correct settings

ELOG V3.1.4-966e3dd